# Distributionally Robust Markov Games with Average Reward

### v2

@@ 7 Connections with Discounted DR-NE
$V^{\pi^{\star}}_{\mathcal{P},i}\geq V^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (9)
where $V^{\pi^{\star}}_{\mathcal{P},i}$ is the $\gamma$-discounted robust value function. A policy $\pi^{\star}$ is an $\epsilon$-Nash Equilibrium, if
$g^{\pi^{\star}}_{\mathcal{P},i}+\epsilon\geq g^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i},\forall i\in\mathcal{N},\forall\pi_{i},$ (10)
@@
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted distributionally robust Markov game by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward distributionally robust Markov game.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted distributionally robust Markov game is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
@@
The importance of this theorem lies in its computational implications. The analysis of $\gamma$-discounted games is often more tractable (kardecs2011discounted), benefiting from properties like the Banach fixed-point theorem which guarantees the existence and uniqueness of value functions. Many reinforcement learning and game theory algorithms, such as value iteration and policy iteration, are naturally formulated for the discounted setting. Our results hence provide a rigorous justification for using these well-understood tools to solve the more challenging average reward problem.
Remark 3.
Part (2) of Theorem 7.1 offers a more practical and constructive guarantee. In practice (like data-driven learning), it may be sufficient to find some approximation of the NE instead of the exact NE. This result therefore assures us that for any desired level of accuracy $\epsilon$, we can simply select a single discount factor $\gamma$ sufficiently close to 1. Any $\epsilon$-Nash Equilibrium found for this $\gamma$-discounted game will be an $\epsilon$-Nash Equilibrium for the average-reward game. This provides a clear and direct pathway for finding near-equilibrium policies for the average-reward criterion. Moreover, we can derive the following result to specify the reduction to discounted DR-MGs.
Lemma 7.1.
Define the robust diameter (auer2008near) of the distributionally robust Markov game by $D=\max_{\mathsf{P}\in\mathcal{P}}\max_{s,s^{\prime}\in\mathcal{S}}\min_{\pi}\mathbb{E}[T(s^{\prime}|s,\pi,\mathsf{P})]$, where $T(s^{\prime}|s,\pi,\mathsf{P})$ is the first (random) time step in which state $s^{\prime}$ is reached when policy $\pi$ is executed with transition kernel $\mathsf{P}$ with initial state $s$. Then the discount factor $\gamma=1-\frac{\epsilon}{D}$ satisfies Part (2) of Theorem 7.1.
The result hence specifies a reduction factor whose Nash Equilibrium is an approximated equilibrium under the average reward criterion. Thus, we can apply existing algorithms for discounted DR-MGs (zhang2020robust) to approximately solve the average reward ones.
8 Numerical Experiments
In this section, we develop numerical experiments to validate the advantages of our formulation and methods. We mainly consider two aspects: the long-term performance and the robustness to environment uncertainty.
8.1 Long-Term Performance
To empirically validate our approach, we perform a series of experiments designed to compare the performance of our robust Nash-Iteration algorithm with average reward against the standard discounted-reward robust Nash-Iteration (zhang2020robust; shi2024sample) under DR-MGs. We aim to show that our algorithm will yield a joint policy that has better long-term performance than those derived from the discounted criterion.
We design a "Structured Random Environment" to create a complex strategic challenge with meaningful trade-offs between short-term gains and long-term consequences. The environment is a DR-MG with 20 states and 5 actions. The state space $S$ is partitioned into two clusters:
  • Prosperous States ($|S_{p}|=5$): A small set of states characterized by a higher mean reward (drawn from $\mathsf{N}(2,1)$).
  • Deprived States ($|S_{d}|=15$): A larger set of states with a lower mean reward (drawn from $\mathsf{N}(-2,1)$),
where $\mathsf{N}(\cdot,\cdot)$ is the Gaussian distribution. The nominal transition kernel $\mathsf{P}_{0}$ is also structured to create persistence within these clusters. Transitions originating from a state in a given cluster are five times more likely to terminate in a state within the same cluster. This design creates a strategic trap: an agent in a prosperous state might be tempted by an action with a high immediate reward that carries a significant risk of transitioning it to the deprived cluster, from which escape is difficult.
We further set the uncertainty set using a Kullback-Leibler (KL) divergence ball around the nominal kernel $\mathsf{P}_{0}$:
$\mathcal{P}^{a}_{s}=\{q\in\Delta(\mathcal{S})\ |\ D_{KL}(q||(\mathsf{P}_{0})^{a}_{s})\leq\theta\}.$
Based on the scale of our environment, the uncertainty budget $\theta$ is set to a modest value of 0.01.
We compare our average-reward Robust Nash-Iteration algorithm (Algorithm 1) and the Discounted-Reward Robust Nash-Iteration. We implement the two algorithms and evaluate their output policies under the robust average-reward criterion. The experimental procedure is as follows:
(1). The optimal average-reward policy, $\pi_{avg}$, is computed using the average-reward Robust Nash-Iteration algorithm. We evaluate the worst-case average reward of Player 1 under the policy, using the robust average reward single-agent RL method in (wang2023robust).
(2). For a range of discount factors $\gamma$ from 0.5 to 0.99, the corresponding optimal discounted policy, $\pi_{disc}(\gamma)$, is computed using the Discounted-Reward Robust Nash-Iteration algorithm. And for each $\pi_{disc}(\gamma)$, we similarly evaluate its worst-case long-term performance.
For both algorithms, we plot the performance of Player 1 as a performance metric. Results are presented in Figure 1 (under different random environments). Policies learned under the discounted criterion consistently underperform the dedicated average-reward policy. This gap is most pronounced for smaller values of $\gamma$, where the myopic nature of the discounted algorithm leads it to fall into the strategic trap of the environment by choosing actions that yield high immediate rewards but lead to the persistent, low-reward "deprived" cluster of states. As $\gamma$ approaches 1, the discounted policy becomes more far-sighted, and its performance converges toward the optimal baseline.
Our results hence provide strong evidence that, for DRMGs under the average reward criterion with long-term performance optimization goals, our robust Nash-Iteration algorithm achieves a better performance, compared to the common practice of approximating this objective with a discounted formulation using $\gamma\approx 1$.
Figure 1: Performance Comparisons of Average and Discounted Robust Nash-Iteration
Figure 2: Performance Comparisons of Robust/Non-Robust Nash-Iteration
8.2 Robustness
We then test the robustness of our method by comparing it with the non-robust average-reward Nash-Iteration algorithm (li2003learning), to test their ability to handle model uncertainty. Using the same environment, we evaluate both learners against a worst-case adversary throughout training to measure their inherent robustness. The results are presented in Figure 2. The policy learned by the robust algorithm consistently and monotonically improves its worst-case performance, rapidly converging to a high, stable value. This is because the algorithm directly optimizes for resilience against the adversary at each step. In contrast, the policy learned by the non-robust algorithm exhibits lower performance when evaluated against a worst-case adversary. Ultimately, it does not have the ability to reliably account for perturbations in the underlying environment, as opposed to its robust counterpart, which aligns with our theoretical results.
These numerical results therefore verify that our method is able to achieve superior long-term performance with increased stability when in the face of model uncertainty in multi-agent systems, validating our theoretical results.
@@ 9 Conclusion
In this paper, we developed a comprehensive framework for distributionally robust Markov games under the average-reward criterion, a setting critical for multi-agent systems requiring long-term reliability in the face of model uncertainty. We established the well-posedness of the robust average-reward Bellman equation, proved the existence of a stationary Nash Equilibrium under both irreducible and weak communicating settings, and proposed the Robust Nash-Iteration algorithm as the first provably convergent method for this game. We also connected average-reward equilibria to their discounted counterparts, and our numerical experiments validated the superior long-term performance and stability of our approach. This work provides a complete theoretical and algorithmic blueprint for designing reliable multi-agent systems that can navigate significant environmental uncertainty over extended horizons.
@@ A.1 Discounted Setting
$V^{\pi}_{\mathcal{P},\gamma}(s)\triangleq\min_{\kappa\in\bigotimes_{t\geq 0}\mathcal{P}}\mathbb{E}_{\pi,\kappa}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|S_{0}=s\right],$ (11)
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\bigotimes_{t\geq 0}\mathcal{P}$.
For robust discounted MDPs, it has been shown that the robust discounted value function is the unique fixed-point of the robust discounted Bellman operator (nilim2004robustness; iyengar2005robust; puterman1994markov):
@@ A.2 Average Reward Setting
We can view equation (16) as the cumulative difference over time $t$ as $t\rightarrow\infty$. Defining $H^{\pi}_{\mathsf{P}_{0}}\triangleq(I-\mathsf{P}_{0}^{\pi}+\mathsf{P}_{*}^{\pi})^{-1}(I-\mathsf{P}_{*}^{\pi})$ as the deviation matrix of $\mathsf{P}_{0}^{\pi}$, then $h_{\mathsf{P}_{0}}^{\pi}=H^{\pi}_{\mathsf{P}_{0}}r_{\pi}$ by (puterman1994markov), thus it holds the following non-robust Bellman equation:
@@
$g_{\mathcal{P}}^{\pi}(s)$ $\triangleq\min_{\kappa\in\bigotimes_{t\geq 0}\mathcal{P}}\liminf_{n\to\infty}\mathbb{E}_{\pi,\kappa}\left[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\right].$ (18)
@@
For the robust Markov game above, no (stationary mixed) Nash equilibrium exists.
@@ Appendix C Irreducible DR-MGs
Consider an arbitrary uncertainty set, $\mathcal{P}$, satisfying Assumption 1, and denote any worst-case transition kernel w.r.t. $g^{\pi}_{\mathsf{P}}$ by $\mathsf{P}_{w}\triangleq\arg\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P}}$. Then the robust Bellman equation:
$V=r^{\pi}-g+\sigma^{\pi}(V)$ (21)
has a solution $(g^{\pi}_{\mathsf{P}_{w}},V^{\pi}_{\mathsf{P}_{w}})$.

### v3

@@ 7 Connections with Discounted DR-NE
$V^{\pi^{\star}}_{\mathcal{P},i}(\rho)\geq V^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i},$ (13)
where $V^{\pi^{\star}}_{\mathcal{P},i}(\rho)$ is the $\gamma$-discounted robust value function. A policy $\pi^{\star}$ is an $\epsilon$-Nash Equilibrium, if
$g^{\pi^{\star}}_{\mathcal{P},i}(\rho)+\epsilon\geq g^{(\pi^{\star}_{-i},\pi_{i})}_{\mathcal{P},i}(\rho),\forall i\in\mathcal{N},\forall\pi_{i},$ (14)
@@
(1). For a sequence of discount factors $\{\gamma_{t}\}\to 1$, denote the Nash Equilibrium of the $\gamma_{t}$-discounted DR-MG by $\pi_{t}$. If $\pi_{t}\to\pi$, then $\pi$ is a Nash Equilibrium of the average-reward DR-MG.
(2). For any $\epsilon$, there exists some discount factor $\gamma$, such that any $\epsilon$-Nash Equilibrium of the $\gamma$-discounted DR-MG is also an $\mathcal{O}(\epsilon)$-Nash Equilibrium under the average reward.
(3). Moreover, define the robust diameter (auer2008near) of the DR-MG by $D=\max_{\mathsf{P}\in\mathcal{P}}\max_{s,s^{\prime}\in\mathcal{S}}\min_{\pi}\mathbb{E}[T(s^{\prime}|s,\pi,\mathsf{P})]$, where $T(s^{\prime}|s,\pi,\mathsf{P})$ is the first (random) time step in which state $s^{\prime}$ is reached when policy $\pi$ is executed with transition kernel $\mathsf{P}$ with initial state $s$. Then the discounted robust NE for $\gamma=1-\frac{\epsilon}{D}$ is an $\epsilon$-NE under average reward.
@@
Our results provide significant computational implications. The analysis of $\gamma$-discounted games is often more tractable (kardecs2011discounted), benefiting from properties like the Banach fixed-point theorem, which guarantees the existence and uniqueness of value functions. Many reinforcement learning and game theory algorithms, such as Q-learning and policy gradient, can be applied for the discounted setting (zhang2020robust). This result therefore assures us that for any desired level of accuracy $\epsilon$, we can simply select a single discount factor $\gamma$ sufficiently close to 1 and solve for this $\gamma$-discounted game.
@@ 8 Conclusion
In this paper, we developed a comprehensive framework for distributionally robust Markov games under the average-reward criterion, a setting critical for multi-agent systems requiring long-term reliability in the face of model uncertainty. We proved the existence of a stationary Nash Equilibrium under both irreducible and weak communicating settings, and proposed two practical algorithms as the first provably convergent method for average reward DR-MGs. We also connected average-reward equilibria to their discounted counterparts. This work provides a complete theoretical and algorithmic blueprint for designing reliable multi-agent systems that can navigate significant environmental uncertainty over extended horizons.
@@ A.1 Discounted Setting
$V^{\pi}_{\mathcal{P},\gamma}(s)\triangleq\min_{\kappa\in\times_{t\geq 0}\mathcal{P}}\mathbb{E}_{\pi,\kappa}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}|S_{0}=s\right],$ (15)
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\times_{t\geq 0}\mathcal{P}$.
For robust discounted MDPs, it has been shown that the robust discounted value function is the unique fixed-point of the robust discounted Bellman operator (nilim2004robustness; iyengar2005robust; puterman2014markov):
@@ A.2 Average Reward Setting
We can view equation (20) as the cumulative difference over time $t$ as $t\rightarrow\infty$. Defining $H^{\pi}_{\mathsf{P}_{0}}\triangleq(I-\mathsf{P}_{0}^{\pi}+\mathsf{P}_{*}^{\pi})^{-1}(I-\mathsf{P}_{*}^{\pi})$ as the deviation matrix of $\mathsf{P}_{0}^{\pi}$, then $h_{\mathsf{P}_{0}}^{\pi}=H^{\pi}_{\mathsf{P}_{0}}r_{\pi}$ by (puterman2014markov), thus it holds the following non-robust Bellman equation:
@@
$g_{\mathcal{P}}^{\pi}(s)$ $\triangleq\min_{\kappa\in\times_{t\geq 0}\mathcal{P}}\liminf_{n\to\infty}\mathbb{E}_{\pi,\kappa}\left[\frac{1}{n}\sum^{n-1}_{t=0}r_{t}|S_{0}=s\right].$ (22)
@@
For the robust Markov game above, no (stationary stochastic) Nash equilibrium exists.
@@ Appendix C Irreducible DR-MGs
In this appendix we work in the single-agent setting and fix a stationary policy $\pi$. The state space $\mathcal{S}$ and action space $\mathcal{A}$ are finite. For each $(s,a)$, the local ambiguity set $\mathcal{P}_{s}^{a}\subset\Delta(\mathcal{S})$ is nonempty, compact and convex, and the overall ambiguity set is the rectangular product
$\mathcal{P}=\times_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}_{s}^{a}.$
For each $P\in\mathcal{P}$, we denote by $P^{\pi}$ the induced Markov kernel on $\mathcal{S}$:
$P^{\pi}(s^{\prime}\mid s)\;=\;\sum_{a\in\mathcal{A}}\pi(a\mid s)P(s^{\prime}\mid s,a).$
Under Assumption 1, $P^{\pi}$ is irreducible for every stationary $\pi$ and for every $P\in\mathcal{P}$.
For fixed $P$ and $\pi$, the average reward (gain) and bias are defined as
$g^{\pi}_{P}(s)=\liminf_{n\to\infty}\mathbb{E}^{\pi,P}\!\left[\frac{1}{n}\sum_{t=0}^{n-1}r_{t}\,\middle|\,S_{0}=s\right],\qquad h^{\pi}_{P}(s)=\mathbb{E}^{\pi,P}\!\left[\sum_{t=0}^{\infty}\bigl(r_{t}-g^{\pi}_{P}\bigr)\,\middle|\,S_{0}=s\right].$
By irreducibility, $g^{\pi}_{P}(s)$ does not depend on $s$, so we write simply $g^{\pi}_{P}$; moreover $(g^{\pi}_{P},h^{\pi}_{P})$ is a solution to the (non-robust) Poisson/Bellman equation (puterman2014markov):
$g^{\pi}_{P}+h^{\pi}_{P}(s)\;=\;r^{\pi}(s)+\sum_{s^{\prime}\in\mathcal{S}}P^{\pi}(s^{\prime}\mid s)\,h^{\pi}_{P}(s^{\prime}),\quad s\in\mathcal{S},$ (25)
unique up to an additive constant on the bias.
We define the robust average reward of $\pi$ as
$g^{\pi}_{\mathcal{P}}\;\triangleq\;\min_{P\in\mathcal{P}}g^{\pi}_{P}.$
Compactness of $\mathcal{P}$ and continuity of $P\mapsto g^{\pi}_{P}$ (see, e.g., Puterman, 2014, Prop. 8.4.6) imply that the minimum is attained: there exists at least one worst-case kernel
$P^{w}\in\arg\min_{P\in\mathcal{P}}g^{\pi}_{P}.$
For each state $s$ and action $a$, define the local robust one-step operator
$\sigma_{s}^{a}(v)\;\triangleq\;\min_{p\in\mathcal{P}_{s}^{a}}p^{\top}v,\qquad v\in\mathbb{R}^{\mathcal{S}},$
and the $\pi$-averaged robust Bellman operator
$\sigma^{\pi}_{s}(v)\;\triangleq\;\min_{P\in\mathcal{P}}\sum_{s^{\prime}\in\mathcal{S}}P^{\pi}(s^{\prime}\mid s)v(s^{\prime})\;=\;\min_{p\in\mathcal{P}_{s}^{\pi}}p^{\top}v,$
where $\mathcal{P}_{s}^{\pi}\triangleq\bigl\{\sum_{a}\pi(a\mid s)\,p^{a}:p^{a}\in\mathcal{P}_{s}^{a}\bigr\}$.
@@
Consider an uncertainty set $\mathcal{P}$ satisfying Assumption 1 and fix a stationary policy $\pi$. Let $P^{w}\in\arg\min_{P\in\mathcal{P}}g^{\pi}_{P}$ be any worst-case kernel, and let $(g^{\pi}_{P^{w}},h^{\pi}_{P^{w}})$ be the corresponding (non-robust) gain and bias, i.e. a solution to (25) with $P=P^{w}$. Then the pair $(g^{\pi}_{\mathcal{P}},h^{\pi}_{P^{w}})$ solves the robust Bellman equation
$g^{\pi}_{\mathcal{P}}+h(s)\;=\;r^{\pi}(s)+\sigma^{\pi}_{s}(h),\qquad s\in\mathcal{S},$ (26)
and is unique up to addition of a constant to $h$. Equivalently, writing $V=h$ and rearranging, $V$ satisfies
$V(s)\;=\;\sum_{a\in\mathcal{A}}\pi(a\mid s)\bigl(r(s,a)-g^{\pi}_{\mathcal{P}}+\sigma_{s}^{a}(V)\bigr),\quad s\in\mathcal{S},$
which is exactly equation (6) in the main text.